# Disorder to Order Transition in 1D Nonreciprocal Cahn-Hilliard Model

### v1

@@ V Conclusions
In this paper, we have presented a systematic study of the disorder to order transition for the 1D NRCH model. In 1D, defect solutions are the sources and sinks of travelling waves. Sources are analogous to the two-dimensional topologically neutral targets, whereas sinks are similar to the disinclination lines where the waves emanated from sources meet. For a given $\alpha$, defects select a unique wave number $k_{\infty}$ that increases monotonically with $\alpha$. Wave number selection predicts a crossover threshold $\alpha_{\times}\sim 0.62$ above which the defect solutions cannot exist.
@@
At small $\alpha$, disordered initial states evolve into multi-defect configurations, where sources and sinks are arranged in an alternating manner. With increasing $\alpha$, overall the defect density increases, peaks at $\alpha\sim 0.2$, and then vanishes at $\alpha_{c}\sim 0.6$, which marks the onset of disorder to order transition. In accordance with this observation, we find that there is no significant average polar order for $\alpha<\alpha_{c}$, but travelling states for $\alpha>\alpha_{c}$ show perfect order in the long time steady state. In a direct contrast to 2D, where the transition occurs at $\alpha_{c}\sim 0.28\ll\alpha_{\times}\sim 0.58$, the transition point in 1D, $\alpha_{c}\sim 0.6$ is in close agreement with the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by wavenumber selection. A closer inspection of the defect dynamics reveals that at certain resonance values of nonreciprocity below $\alpha_{c}$, $\rho_{D}$ shows sharp minima, which is a consequence of the fact that for these values of $\alpha$, defect merger events continue to occur even at very long times. Additionally, the average inter-defect separation rises sharply for these $\alpha_{c}$, which was also observed in the study of pairwise defect interactions in 2D [4].
@@
Our numerical simulations with N-BC and D-BC, which are inconsistent with the travelling waves, show different flavours of the disorder-order transition. While the phenomenology for the N-BC is similar to the P-BC for $\alpha<\alpha_{c}$, D-BC significantly affects the defect dynamics and below $\alpha_{c}$, we find configurations with one, many, or no defects at all. Above $\alpha_{c}$, N-BC and D-BC share similar phenomenology, which is distinct from the P-BC. For nonreciprocity slightly above the threshold, we find intermittent, fluctuating polar order. For larger values of $\alpha$, the system partitions into two subdomains which sustain over time.
@@
To conclude, our study shows that for the P-BC NRCH model shows similar phenomenological behaviour in both 1D and 2D, with a crucial difference that the disorder-to-order transition occurs at the crossover value predicted by wave number selection in 1D. We further show that boundary conditions can significantly affect the phenomenology. These findings call for future studies to further investigate the defect dynamics and defect interactions for conserved systems with nonreciprocal interactions.

### v2

@@ V Conclusions
We have presented a systematic study of the disorder-to-order transition for the 1D NRCH model. In 1D, defect solutions are the sources and sinks of travelling waves. Sources are analogous to the two-dimensional topologically neutral targets, whereas sinks are similar to the disclination lines where the waves emanated from sources meet. For a given $\alpha$, defects select a unique wave number $k_{\infty}$ that increases monotonically with $\alpha$. Wave number selection predicts a crossover threshold $\alpha_{\times}\sim 0.62$ above which the defect solutions cannot exist.
@@
At small $\alpha$, disordered initial states evolve into multi-defect configurations, where sources and sinks are arranged in an alternating manner. With increasing $\alpha$, overall the defect density increases, peaks at $\alpha\sim 0.2$, and then vanishes at $\alpha_{c}\sim 0.6$, which marks the onset of disorder-to-order transition. In accordance with this observation, we find that there is no significant average polar order for $\alpha<\alpha_{c}$, but travelling states for $\alpha>\alpha_{c}$ show perfect order in the long time steady state. In a direct contrast to 2D, where the transition occurs at $\alpha_{c}\sim 0.28\ll\alpha_{\times}\sim 0.58$, the transition point in 1D, $\alpha_{c}\sim 0.6$ is in close agreement with the crossover threshold $\alpha_{\times}\sim 0.62$ predicted by wavenumber selection. A closer inspection of the defect dynamics reveals that at certain resonance values of nonreciprocity below $\alpha_{c}$, $\rho_{D}$ shows sharp minima, which is a consequence of the fact that for these values of $\alpha$, defect merger events continue to occur even at very long times. Additionally, the average inter-defect separation rises sharply for these $\alpha$ values, which was also observed in the study of pairwise defect interactions in 2D [4].
@@
Our numerical simulations with N-BC and D-BC, which are inconsistent with the travelling waves, show a different flavor of the disorder-order transition. The defect dynamics at $\alpha<\alpha_{c}$ is similar to the P-BC, but the boundary conditions significantly affects the phenomenology above the transition point. Slightly above the threshold, we find intermittent patches of polar order. For larger values of $\alpha$, the system partitions into two subdomains which sustain over time.
@@
To conclude, our study shows that for the P-BC NRCH model shows similar phenomenological behaviour in both 1D and 2D, with a crucial difference that the disorder-to-order transition occurs at the crossover value predicted by wave number selection in 1D. We further show that boundary conditions can significantly affect the phenomenology. These findings call for future studies to further investigate the defect dynamics and defect interactions for conserved systems with non-reciprocal interactions. In particular, it will be interesting to explore what determines the value of $\alpha_{c}$ in 1D and 2D and to find out the mechanism that drives the resonances for certain values of $\alpha$.